OpenStudy (anonymous):

3x + y = -4 5x - y = -8

OpenStudy (anonymous):

Please Look at your other question as i asked a team to help you

OpenStudy (anonymous):

This one has different symbols but the same numbers and variables. But thanks!

OpenStudy (anonymous):

Oh one moment i can call them here lol

OpenStudy (anonymous):

Asked mods (Smartys here )

OpenStudy (amistre64):

seems like a trap :)

OpenStudy (amistre64):

what do we get when we add the equations together?

OpenStudy (anonymous):

Thanks i am heading to my Computer Science Category

OpenStudy (amistre64):

3x + y = -4 + 5x - y = -8 -------------- 8x + 0 = -12

OpenStudy (amistre64):

ack!!, now i have to edit that lol

OpenStudy (amistre64):

so, we know that "x" at least has to = -12/8 use that to determine a good value for "y"

OpenStudy (anonymous):

I have x =3.42

OpenStudy (amistre64):

sounds good to me as long as it fits :)

OpenStudy (amistre64):

-36 + 8y = -32 8y = 4 y = 4/8 maybe

OpenStudy (amistre64):

i dont think x = 3.42 tho, seems a but off

OpenStudy (amistre64):

-12/8 = -3/2 = -1.5 in my book